# Privacy-Preserving Collaborative Split Learning Framework for Smart Grid Load Forecasting

### v2

@@ IV-F Discussion
In Section IV-E, we analyzed the extent of privacy leakage arising from the sharing of clients’ activations using MINE. In Fig. 8, we showed that even without the added noise, the Split-1 activations have significantly low MI w.r.t. the inputs due to their non-linear and complex relationship induced by the Split-1 model. Moreover, based on the estimated MI, a client can decide whether to forward the current batch activations to GS or not to mitigate privacy leakage. We further analyze the effects of introducing differential privacy as an additional layer of security on the model’s performance. In Fig. 10, we see that with a moderate privacy budget of $\epsilon=5.0$, the models performed similarly to the non-DP case and saw performance degradation only when $\epsilon\leq 2.5$, leading to strong privacy. With a trained model, the electricity service provider can perform individual-level predictions (requiring respective clients’ involvement) and neighbourhood-level predictions using the cumulative load trend from the grid station servicing the neighbourhood. The LSTM-based FL models [22, 19] proposed so far were able to make 1h to 24h ahead predictions; our transformer-based model can perform accurate predictions from 24h - 720h [16], although our experiments were limited to 96h ahead predictions only.
@@ V Conclusion
In this article, we propose a split learning framework to train a DL time series prediction model using the client smart meter data without compromising individual clients’ privacy. Our proposed SL frameworks use the smart meters to perform a forward pass through the Split-1 network only, while the rest of the training is relegated to GS and SP entities. This ensures that the main functionalities of the smart meter remain unhindered. Once trained, the entire model is retained by the energy provider and can be used to perform load predictions for a single smart meter or the entire neighbourhood. The experimental results have shown the performance of the trained models to be better or on par with a centrally trained model. To analyze the extent of information leakage through the Split-1 network, we used mutual information neural estimation to approximate the MI between the input and output of the Split-1 network. The analysis showed that the MI leakage through the Split-1 network is limited. Furthermore, as an added layer of security, we analyzed the addition of $\epsilon$-DP and $(\epsilon,\delta)$-DP to our framework under multiple privacy budgets. We found that the models performed similarly to the non-DP case under a medium privacy budget while observing low-performance degradation under relatively low privacy budgets.

### v3

@@ IV-H Discussion
In Section IV-E, we analyzed the extent of privacy leakage arising from the sharing of clients’ activations using MINE. In Fig. 7, we showed that even without the added noise, the Split-1 activations have significantly low MI w.r.t. the inputs due to their non-linear and complex relationship induced by the Split-1 model. Moreover, based on the estimated MI, a client can decide whether to forward the current batch activations to GS or not to mitigate privacy leakage. We further analyze the effects of introducing differential privacy as an additional layer of security on the model’s performance. In Fig. 9, we see that with a moderate privacy budget of $\epsilon=5.0$, the models performed similarly to the non-DP case and saw performance degradation only when $\epsilon\leq 2.5$, leading to strong privacy. With a trained model, the electricity service provider can perform individual-level predictions (requiring respective clients’ involvement) and neighbourhood-level predictions using the cumulative load trend from the grid station servicing the neighbourhood.
In Section IV-F we evaluated the SplitGlobal framework under FSHA [40] in both non-DP and DP-enabled settings. In the non-DP case, the attacker’s reconstruction achieved an MSE of 0.241, while enabling DP increased the reconstruction MSE to 0.328, indicating a clear improvement in privacy protection through DP regularization. Finally, in Section IV-G, our overhead analysis shows that the client-side in the Split learning framework contributes only approx. $5$% of the total computational load and $4.55$% of the communication load per training round. Additionally, under reasonable hardware assumptions, the client accounts for approximately $21$% of the end-to-end latency and similar proportions in energy consumption, highlighting the lightweight nature of the client-side operations.
Our proposed split learning framework is well-suited for integration into modern power systems, particularly within infrastructures like Advanced Metering Infrastructure (AMI) and Energy Management Systems (EMS). The Split-1 module can be embedded in smart meters, while the GS and SP components can be deployed at the utility’s backend, leveraging existing EMS or Distribution Management System (DMS) facilities. This decentralized architecture supports scalable and privacy-preserving model training across diverse consumer endpoints. However, practical deployment may face challenges such as communication latency, device heterogeneity, synchronization issues, and secure data exchange. Addressing these challenges will be essential to fully realize the framework’s potential in real-world settings, and we identify these areas as important directions for future work.
@@ V Conclusion
In this article, we propose a split learning framework to train a DL time series prediction model using the client smart meter data without compromising individual clients’ privacy. Our proposed SL frameworks use the smart meters to perform a forward pass through the Split-1 network only, while the rest of the training is relegated to GS and SP entities. This ensures that the smart meter’s main functionalities remain unhindered. Once trained, the energy provider retains the entire model, which can be used to perform load predictions for a single smart meter or the entire neighbourhood. The experimental results have shown that the performance of the trained models is better or on par with that of a centrally trained model. To analyze the extent of information leakage through the Split-1 network, we used mutual information neural estimation to approximate the MI between the input and output of the Split-1 network. The analysis showed that the MI leakage through the Split-1 network is limited. Furthermore, as an added layer of security, we analyzed the addition of $\epsilon$-DP and $(\epsilon,\delta)$-DP to our framework under multiple privacy budgets. We found that the models performed similarly to the non-DP case under a medium privacy budget while observing low-performance degradation under relatively low privacy budgets.
Appendix A Comprehensive Overhead Analysis of the Proposed Split Learning Framework
To provide a detailed overhead analysis of the proposed framework, we refer to the Dual Split FEDformer model illustrated in Fig. 2. The relevant model parameters and their values are listed in Table VI.
In the discussion that follows, we assume a batch size of $1$, which aligns with the per-client data processing during a single training epoch. This choice also reflects the trade-off between training speed and the computational and communication capabilities available at the client (smart meter) side. The analysis below is focused on a single round of model training.
TABLE VI: Model Parameters
Symbol Description Value
$C$ Clients 10
$B$ Batch Size 1
$N$ FFT Size 128
$D$ Model dimension 512
$D_{ff}$ FFN Dimension 2048
$M$ Modes of FFT kept 64
$L_{e}$ Encoder Length 96
$L_{d}$ Decoder Length (1.5 x $L_{e}$) 144
$O$ Prediction Horizon 96
A-A Communication Overhead:
We break down the communication overhead into forward and backward pass components:
  • Forward Pass: Each smart meter handles the forward pass through the Split-1 block, computes the final loss, and initiates gradient backpropagation. The set of Split-1 activations sent from each client to the GS, denoted $\textbf{A}_{k,b}^{GS}$ in Algorithm 1, includes the output tensors from the:
    • Split-1 Encoder: $X_{enc}^{S_{1}}\in\mathbb{R}^{B\times L_{e}\times D}$
    • Split-1 Decoder: $S_{dec}^{S_{1}}\in\mathbb{R}^{B\times L_{d}\times D}$
    • Trend component from Split-1 Decoder: $trend_{enc}^{S_{1}}\in\mathbb{R}^{B\times L_{d}\times D}$.
    Using the model parameters in Table VI and assuming Float32 precision, the payload size per client for these activations is approximately $6.3$ MB. The GS aggregates the activations from all $C=10$ clients and transmits them to the SP, resulting in a total payload of approximately $63$ MB.
    For loss computation, the SP returns a small output tensor of size $B\times O$ (roughly $3$ KB per client) back to the respective smart meters via the GS, this is negligible in terms of communication overhead.
  • Backward Pass: The client initiates backpropagation by sending the gradient of the loss (approx. $3$ KB) to the SP via the GS. The SP continues backpropagation, computes gradients for Split-1, and sends them (approx. $63$ MB) to the GS. The GS completes the backpropagation and updates the Split-1 model, while the SP updates the Split-2 model.
    In total, the communication volume during a single model update iteration is: $6.3$ MB (client $\rightarrow$ GS) + $63$ MB (GS $\rightarrow$ SP) + $63$ MB (SP $\rightarrow$ GS) = $132.3$ MB
    Notably, only 4.55% of the total communication is from the client to the GS; the majority occurs between GS and SP. Please note that these calculations exclude transmission protocol overhead.
A-B Computational Overhead:
Estimating the total computational cost for a single round of model training is inherently complex due to backend optimizations in frameworks such as PyTorch. However, we can approximate the number of operations by counting the number of multiplications (ignoring additions) performed within key modules of the FEDformer model. These estimates are based on the methodology provided in Section 4.2 of [16].
The most computationally intensive modules are the Frequency Enhanced Block (FEB), Frequency Enhanced Attention (FEA), and Feed Forward Network (FFN). The Series Decomposition Block, which uses lightweight convolutional operations, is negligible in comparison and is omitted from the analysis. All notations follow those in Table VI.
Module-wise multiplication count:
  • FEB:
    • Encoder: $\approx\left[(L_{e}+M)D^{2}+2NL_{e}D\right]$,
    • Decoder: $\approx\left[(L_{d}+M)D^{2}+2NL_{d}D\right]$.
  • FEA: $\approx\left[(2L_{e}+L_{d})D^{2}+2(L_{e}+L_{d})ND+2M^{2}D\right]$.
  • FFN:
    • Encoder: $\approx 2L_{e}DD_{ff}$,
    • Decoder: $\approx 2L_{d}DD_{ff}$
Using these expressions, we now compute the computational load distribution between the client and the GS-SP infrastructure for a single forward pass.
  • The Client Split-1: As shown in Fig. 2, Split-1 comprises two FEB blocks (one in Encoder, one in Decoder), and a projection via matrix $W_{1}\in\mathbb{R}^{D\times Z}$. For our univariate case ($Z=1$), the projection involves $L_{d}D$ multiplications. Substituting $L_{d}=1.5L_{e}$ and using $L_{e}=L$, the total number of multiplications becomes:
    $\textrm{Ops}_{\textrm{Client}}=(2.5L+2M)D^{2}+(5N+1.5)LD$ (8)
    which evaluates to approximately $1.28\times 10^{8}$ multiplications.
  • The SP Split-2: The SP side consists of:
    • Encoder Layers: 2 FFN modules and 1 FEB
    • Decoder Layer: 1 FEA block, 1 FFN block, and 3 projection operations
    The total multiplication count is:
    $\textrm{Ops}_{\textrm{SP}}=(4D_{ff}+7N+4.5)LD+(4.5L+M)D^{2}+2M^{2}D+3LDD_{ff}$ (9)
    which evaluates to approximately $8.83\times 10^{8}$ multiplications. To summarize, in a forward pass through the network, the client performs approx. 12.66% of the entire operations.
Once the forward pass is complete, the gradient backpropagation through Split-2 is handled by the SP, and through Split-1 by the GS. The client only participates in computing the loss and initiating the backward pass. Typically, the backward pass incurs 23×2–3$\times$2 – 3 × the cost of the forward pass due to additional gradient calculations, activation derivatives, and layer-specific operations (e.g., dropout, batch normalization). Taking a conservative estimate of 2.5×2.5$\times$2.5 × the forward pass cost, the client’s overall contribution in one full training iteration (forward + backward) comes out to be approximately 5% of the total operations.
To summarize, In a single training round, the smart meter (client) is responsible for:
  • $\approx 4.55$% of the total communication overhead,
  • $\approx 5$% of the total computational overhead.
A-C Latency Overhead
As discussed earlier, the client handles approximately $12.66$% of the total operations during the forward pass. Similarly, during the backward pass, the GS processes an equivalent portion, with the remainder executed by the SP. Therefore, the end-to-end latency is predominantly influenced by the utility provider’s infrastructure (GS and SP entities).
This latency can be effectively minimized by deploying high-performance hardware, particularly CUDA-enabled GPUs, which offer significant acceleration over even high-end CPUs for matrix-heavy operations typical in deep learning. As reported in Section IV-A of the manuscript, a single training epoch takes approximately $20$ minutes for both SplitGlobal and SplitPersonal frameworks, and $40$ minutes for the Central model.
It’s important to note that our implementation was conducted on a single machine, without deploying actual smart meters and separate GS-SP hardware. Thus, while we cannot provide a real-world latency measurement, the breakdown presented above offers a framework for estimating latency under practical deployment scenarios.
To illustrate this, consider the following assumptions:
  • The GS-SP setup performs multiplication operations $5\texttimes$ faster than the client device.
  • The GS-SP together are responsible for $95$% of the operations and can complete these in $10$ seconds.
Under these assumptions, the entire round of model update (if only performed by GS-SP) would take $10/0.95\approx 10.53$ seconds. Consequently, the client’s $5$% share would take $0.053\times 5\approx 2.64$ seconds.
Hence, the total latency per round would be approximately $12.64$ seconds, with the client-side accounting for around $21$% of the total latency.
A-D Energy Overhead
Building on the previous discussions, and assuming identical hardware architecture across the client, GS, and SP systems, the client would be responsible for only around $5$% of the total energy consumption during a single model training round, while the remaining $95$% would be incurred by the GS-SP infrastructure.
In our system model, the utility provider manages both the GS and SP components, allowing us to consider the architecture as being divided into two principal entities: the provider and the client. Within this setting, offloading just $5$% of the total computation to the client achieves a significant accuracy gain. As shown in Table II of the revised manuscript, the SplitPersonal model achieves a mean prediction MSE of $0.251$, improving upon the central model’s MSE of $0.308$, which translates to an $18.51$% performance improvement.
If we temporarily disregard the additional communication and orchestration overhead introduced by the split learning framework and assume uniform energy usage per operation across devices, this results in an $18.51$% gain in performance at effectively no extra energy cost to the client. While this is an idealized scenario, it highlights the efficiency potential of the proposed architecture, even when only a minimal portion of the workload is handled by the client device. It is also worth noting that in practical deployments, client devices such as smart meters may be equipped with energy-efficient but lower-performance hardware. In such cases, even though the latency of the client-side computation may increase, the overall energy consumption per operation could be significantly lower compared to high-performance GS-SP infrastructure. This further supports the viability of our split learning approach, as the computationally light client-side load (only around $5$%) could be executed with minimal energy footprint, without compromising model performance.
[Uncaptioned image] Asif Iqbal received the BS degree in Telecommunication Engineering from NUCES-FAST, Peshawar, Pakistan, MS degree in Wireless Communications from LTH, Lunds University, Sweden, and Ph.D in Electrical & Electronics Engineering from The University of Melbourne, Melbourne, Australia in 2008, 2011, and 2019 respectively. He is currently working as a Research Fellow with the Department of Electrical & Computer Engineering at the National University of Singapore, Singapore. Dr. Iqbal previously served on the faculty of NUCES-FAST, Peshawar, Pakistan as an Assistant Professor. His research interests include signal processing, deep learning, sparse signal representations, and privacy-preserving machine learning.
[Uncaptioned image] Prosanta Gope (Senior Member, IEEE) is currently working as an Associate Professor in the Department of Computer Science (Cyber Security) at the University of Sheffield, UK. Dr Gope served as a Research Fellow in the Department of Computer Science at the National University of Singapore (NUS). Primarily driven by tackling challenging real-world security problems, he has expertise in Lightweight Authentication, Authenticated Encryption, 5G and Next Generation Communication Security, Privacy-Preserving Machine Learning, Security in the Internet of Things, Smart-Grid Security, PUF-based security system and IoT Hardware. He has authored more than 100 peer-reviewed articles in several reputable international journals and conferences and has four filed patents. Several of his papers have been published in high-impact journals (such as IEEE TIFS, IEEE TDSC, and IEEE/ACM TON), and prominent security conferences (such as IEEE S&P, ACM CCS, IEEE Computer Security Foundations Symposium (CSF), Privacy Enhancing Technologies Symposium (PETS), ESORICS, Euro S&P, IEEE TrustCom, IEEE HoST, etc.) Dr Gope has been a TPC member and Co-Chair in several reputable international conferences, including PETS, ESORICS, IEEE TrustCom, IEEE GLOBECOM (Security Track), and ARES. He currently serves as an Associate Editor of the IEEE Transactions on Dependable and Secure Computing, IEEE Transactions on Information & Forensics Security, IEEE Transactions on Services Computing, IEEE Systems Journal, and the Journal of Information Security and Applications (Elsevier). His research has been funded by EPSRC, Innovate UK, and the Royal Society.
[Uncaptioned image] Biplab Sikdar (S’98-M’02-SM’09) received the B.Tech. Degree in Electronics and Communication Engineering from North Eastern Hill University, Shillong, India, in 1996, and the M.Tech. Degree in electrical engineering from the Indian Institute of Technology Kanpur, Kanpur, India, in 1998, and the Ph.D. degree in electrical engineering from the Rensselaer Polytechnic Institute, Troy, NY, USA, in 2001. He was a faculty member at the Rensselaer Polytechnic Institute from 2001 to 2013 and was an assistant professor and Associate Professor. He is currently a Professor and Head of the Department of Electrical and Computer Engineering at the National University of Singapore. He also serves as the Director of the Cisco-NUS Corporate Research Laboratory. His current research interests include wireless networks and security for the Internet of Things and cyber-physical systems. He has served as an Associate Editor for the IEEE Transactions on Communications, IEEE Transactions on Mobile Computing, IEEE Internet of Things Journal and IEEE Open Journal of Vehicular Technology.